OpenStudy (anonymous):

wha are the zeros of f(x)=x^3+8x^2+17x+10? a: -2,1,5 b: 2,-1,-5 c: -2,1,-5 D:-2,-1,-5

OpenStudy (anonymous):

\[x^3+8x^2+17x+10=x^3+x^2+7x^2+7x+10x+10 \\ =x^2(x+1)+7x(x+1)+10(x+1) \\ =(x+1)(x2+7x+10)=(x+1)(x+2)(x+5) \]

OpenStudy (anonymous):

so D is the right answer

OpenStudy (anonymous):

can u help me with some others?

OpenStudy (anonymous):

factoring is a commen way to solve equations like this

OpenStudy (anonymous):

what polynomial has roots of -5,-2 and 4? a: x^3-x^2-22x+40 b:x^3+x^2-22x-40 c:x^3+3x^2-18x-40 d:x^3-3x2-18x+40

OpenStudy (anonymous):

we can write (x+5)(x+2)(x-4)=0

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

am i clear?

OpenStudy (anonymous):

\[(x+5)(x+2)(x-4)=(x^2+7x+10)(x-4) \\ = x^3+7x^2+10x-4x^2-28x-40 \\ =x^3+3x^2-18x-40\]

OpenStudy (anonymous):

C is the right one

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

your welcome

OpenStudy (aliahmed):

d is the answer